OpenStudy (anonymous):

Quick question about L'Hospital's rule: Do you just take the derivative once, or keep taking it the derivative until you get a non-indeterminate form?

OpenStudy (anonymous):

Take for instance the following limit: \[L=\lim_{x\to\infty}\frac{x^2}{e^x}\] Directly substituting yields the indeterminate form \(\dfrac{\infty}{\infty}\). Suppose you apply L'Hopital's rule once: \[L=\lim_{x\to\infty}\frac{2x}{e^x}=\frac{\infty}{\infty}\] Given another indeterminate form, you're allowed to apply the rule again: \[L=\lim_{x\to\infty}\frac{2}{e^x}=0\] So, to answer your question, yes: you may apply L'Hopital's rule as many times as necessary. However, there are some limits for which repeated applications of L'Hopital's rule won't work. For example, \[\lim_{x\to\infty}\frac{\sqrt{x+2}}{\sqrt{x-1}}=\frac{\infty}{\infty}\\ \lim_{x\to\infty}\frac{\frac{1}{2\sqrt{x+2}}}{\frac{1}{2\sqrt{x-1}}}=\lim_{x\to\infty}\frac{\sqrt{x-1}}{\sqrt{x+2}}\] When this kind of situation arise, you'll have to look for other methods to compute the limit.

OpenStudy (anonymous):

Thanks so much, I get it now

OpenStudy (anonymous):

You're welcome